Grade 12th passPhysical Chemistry

pressure would be..?
if we decrease the volume by 45%at constant temperature

manisha
7 Years agoGrade 12th pass

1 Answer

Sahani Kumar
7 Years ago
Initial volume = V1
 initial pressure = P2
V2=(V1 –( 45*V1/100))=(55 *V1)/100
P2 =?
 We know that at constant temperature 
P1 * V1=P2 * V2  
on putting the value of V2 we get
P1*V1=P2*(55V1/100)
On solving this equation we get 
P2=( 100 P1/55)
P2=1.81P1
 Regards 
Askiitian member 
If you liked the answer please encourage us by approving the answer.